OpenStudy (anonymous):

How many 3.4kg bags of oranges can be made from a load with a mass of 17/20 tonnes?

OpenStudy (anonymous):

1 ton=1000 kg so 17/20 tons =17/20*1000= 850 kg and 850/3.4= 250 bags